OpenStudy (anonymous):

derive the equation of the parabola with a focus at (-7,5) and a directrix of y=-11

OpenStudy (campbell_st):

lots of ways to do this question  you can use the distance formula. P is a point (x, y) on the parabola and then use the fact the the distance from the focus to P is the same as the distance from P to a point on the directrix D so \[\sqrt{x + 7)^2 + (y - 5)^2} = \sqrt{(x - x)^2 + (y +11)^2}\] the alternative is to find the focal distance = a this the the distance from the vertex to the focus.. the distance from the focus to the directrix is 2a so 2a = 16 then the vertex is a units below the focus on the same line of symmetry x = -7 so the vertex is (-7, 5 - a) then substitute the values into the standard form \[(x - h)^2 = 4a(y - k)\] where (h, k) is the focus and a is the focal length. hope it helps